# Robust Preference for Dynamical Dark Energy in DESI BAO and SN Measurements

### v1

@@ Abstract
I Introduction
One of the most undoubtedly fascinating and unforeseen discoveries of the past three decades is that the Universe is undergoing an accelerated phase of expansion. This was first argued in 1998 through observations of distant Type Ia Supernovae Riess et al. (1998); Perlmutter et al. (1999), and has since been corroborated by a wide variety of other probes Riess et al. (2001); Tegmark et al. (2004); Scranton et al. (2003); Tonry et al. (2003); Knop et al. (2003); Seljak et al. (2005); Feng et al. (2005); Riess et al. (2004); Astier et al. (2006); Eisenstein et al. (2005, 2007); Tegmark et al. (2006); Sahni and Starobinsky (2006); Wood-Vasey et al. (2007); Vikhlinin et al. (2009); Stern et al. (2010); Sherwin et al. (2011); Bennett et al. (2013); Hinshaw et al. (2013); Dawson et al. (2013); de Jong et al. (2013); Anderson et al. (2014); Weinberg et al. (2013); Beutler et al. (2014); Delubac et al. (2015); Betoule et al. (2014); Aubourg et al. (2015); Ross et al. (2015); Moresco et al. (2016a, b); Rubin and Hayden (2016); Alam et al. (2017); Abbott et al. (2016); Haridasu et al. (2017); Troxel et al. (2018); Scolnic et al. (2018); Aghanim et al. (2020a, b); Gómez-Valent (2019); Yang and Gong (2020); Choi et al. (2020); Aiola et al. (2020); Alam et al. (2021); Nadathur et al. (2020); Rose et al. (2020); Di Valentino et al. (2020a); Alam et al. (2021); Asgari et al. (2021); Tröster et al. (2021); Dutcher et al. (2021); Abbott et al. (2022); Moresco et al. (2022); Abbott et al. (2023a); Brout et al. (2022); Madhavacheril et al. (2024); Abbott et al. (2023b); Adame et al. (2024a); Lodha et al. (2024); Abbott et al. (2024); Sánchez et al. (2024); Vincenzi et al. (2024). [^1] For a few caveats, objections, and discussions surrounding this conclusion raised over the years, see Refs. Buchert (2000, 2001, 2008); Hunt and Sarkar (2010); Nielsen et al. (2016); Tutusaus et al. (2017); Dam et al. (2017); Colin et al. (2019); Desgrange et al. (2019); Koksbang (2019a, b); Heinesen (2023).
@@ I Introduction
Since all known forces and components in nature would decelerate the expansion rate of the Universe, acceleration itself requires a physical mechanism beyond the Standard Model of fundamental interactions, able to counteract deceleration, inducing instead a Dark Energy (DE) phase where the dynamics is characterized by negative pressure with an effective Equation of State (EoS) $w<-1/3$ Sahni and Starobinsky (2000); Carroll (2001); Peebles and Ratra (2003); Padmanabhan (2003); Copeland et al. (2006); Caldwell and Kamionkowski (2009); Li et al. (2011); Martin (2012).
@@
In the theoretical framework described by the standard $\Lambda$CDM model of cosmology, DE is parametrized by a positive cosmological constant term ($\Lambda$) in the Einstein equation with its energy density comprising the vast majority of the Universe’s energy budget. Despite its apparent simplicity, this interpretation is not free from conceptual problems and limitations Weinberg (1989); Krauss and Turner (1995); Sahni and Starobinsky (2000); Weinberg (2000); Padmanabhan (2003); Sahni (2002); Yokoyama (2003); Nobbenhuis (2006); Burgess (2015); Joyce et al. (2015); Bull et al. (2016); Wang et al. (2016a). Foremost, plugging a positive cosmological constant component into the gravitational field equations described by General Relativity (GR) implies living in an asymptotically de Sitter universe, which seems to contrast with several theories/models of quantum gravity proposing instead an asymptotically anti-de Sitter universe Brustein and Steinhardt (1993); Witten (2000); Kachru et al. (2003); Polchinski (2006); Danielsson and Van Riet (2018). Secondly, it seems quite natural to question what made us so lucky to live precisely in the cosmic epoch when such a constant component came to be not only relevant but even dominant compared to other contributions such as matter, altering the expansion history of the Universe so prominently as to allow us to become easily aware of its existence and implications Zlatev et al. (1999); Pavon and Zimdahl (2005); Martin (2012); Velten et al. (2014). Finally, and most importantly, when it comes to the physical interpretations, anything that contributes to the energy density of the vacuum behaves akin to a cosmological constant, summing up within the energy-momentum tensor as $T_{\mu\nu}\propto g_{\mu\nu}\rho$ due to Lorentz invariance. Based on standard quantum field theory calculations, one would expect a zero-point energy density contribution of $\rho_{\rm vac}$, which, depending on the ultraviolet wavelength cutoff scale, is found to be somewhere between $10^{50}$ to $10^{120}$ orders of magnitude larger than what is inferred by cosmological data Weinberg (1989). This leads to one of the biggest disagreements between theoretical predictions and observations, arguably requiring a level of fine-tuning that appears to be well beyond what any current theories can realistically explain Dolgov (1997); Weinberg (1989); Straumann (1999); Padmanabhan (2003); Sola (2013). As a result, from a theoretical perspective, the nature of DE remains one of the biggest puzzles in modern physics, sustaining significant research interest in the high-energy physics community. [^2] This diffuse interest is reflected in the wide range of models – both within and beyond the standard cosmological constant – that have been proposed over the years. These include, for example, new (ultra)light fields and modifications to gravity. With no claims to completeness see, e.g., Refs. Amendola (2000); Kamenshchik et al. (2001); Capozziello (2002); Bento et al. (2002); Mangano et al. (2003); Farrar and Peebles (2004); Khoury and Weltman (2004); Li (2004); Amendola et al. (2007); Hu and Sawicki (2007); Cognola et al. (2008); Nojiri et al. (2011); Zhang and Sasaki (2012); Rinaldi (2015a); Luongo and Quevedo (2014); Rinaldi (2015b); De Felice et al. (2016); Wang et al. (2016a); Josset et al. (2017); Burrage and Sakstein (2016); Sebastiani et al. (2017); Nojiri et al. (2017); Burrage and Sakstein (2018); Capozziello et al. (2018); Benisty and Guendelman (2018); Casalino et al. (2018); Yang et al. (2018a); Saridakis et al. (2018); Visinelli and Vagnozzi (2019); Langlois (2019); Benisty et al. (2019); Boshkayev et al. (2019); Heckman et al. (2019); D’Agostino (2019); Mukhopadhyay et al. (2020a); Mukhopadhyay and Majumdar (2019); Mukhopadhyay et al. (2020b); Vagnozzi et al. (2020); Akarsu et al. (2020); Saridakis (2020); Ruchika et al. (2023); Odintsov et al. (2020a, b); Oikonomou (2021a, b); Vagnozzi et al. (2021a); Solanki et al. (2021); Saridakis (2021); Arora et al. (2022); Capozziello et al. (2022); Narawade et al. (2022); D’Agostino et al. (2022); Oikonomou and Giannakoudi (2022); Belfiglio et al. (2023); Luciano and Giné (2023); Kadam et al. (2023); Ong (2023); Bernui et al. (2023); Luciano (2023); Giani and Piattella (2023); Belfiglio et al. (2024); Frion et al. (2023); Adil et al. (2023); Halder et al. (2024).
@@
Despite these remarkable achievements, it is no exaggeration to say that at the time of writing, observations are inconclusive about the physical nature of DE. While too far-away deviations from the canonical $\Lambda$CDM model appear severely constrained Gariazzo et al. (2024), several alternative theoretical frameworks and phenomenological avenues featuring new physics in the dark sector of the model remain at the very least plausible.
@@
Trying to summarize an otherwise very articulated debate, we can adhere to Occam’s razor principle and start considering one of the simplest hypotheses beyond the cosmological constant. This involves assuming that DE can be modeled as a generic fluid with a constant EoS, $w_{0}$. By leaving $w_{0}$ as a free parameter in the theoretical model, cosmological observations can constrain deviations from the cosmological constant value ($w_{0}=-1$). [^4] See, e.g., Refs Bean and Melchiorri (2002); Hannestad and Mortsell (2002); Said et al. (2013); Shafer and Huterer (2014); Zhang (2016); Moresco et al. (2016b); Xu and Zhang (2016); Wang et al. (2016b); Vagnozzi et al. (2017); Zhang (2017); Feng et al. (2018); Scolnic et al. (2018); Wang (2018); Sprenger et al. (2019); Poulin et al. (2018); Roy Choudhury and Naskar (2019); Abbott et al. (2019); Aghanim et al. (2020b); Wang (2021); Roy Choudhury and Hannestad (2020); Alam et al. (2021); Chudaykin et al. (2021); D’Amico et al. (2021); Vagnozzi et al. (2021b); Yang et al. (2021b); Di Valentino (2021); Abbott et al. (2023a); Tröster et al. (2021); Brieden et al. (2022); Grillo et al. (2020); Cao et al. (2022); Zhang et al. (2021); Colgáin et al. (2021); Teng et al. (2021); Krishnan et al. (2021); Nunes and Vagnozzi (2021); Bernardo et al. (2022); Vagnozzi et al. (2022); Bargiacchi et al. (2022); Moresco et al. (2022); Semenaite et al. (2023); Carrilho et al. (2023); Wang (2022); Koussour et al. (2023); Bernardo et al. (2023); Narawade and Mishra (2023); Hou et al. (2023); Escamilla et al. (2024a); Kumar et al. (2023); Bhagat et al. (2023); Mussatayeva et al. (2023); Hogg (2024); Adame et al. (2024b) for recent and not-so-recent discussions and constraints on the DE EoS from a variety of astrophysical and cosmological probes. In this regard, focusing exclusively on the Planck CMB data one might speculate about a preference for a phantom-like DE component ($w_{0}<-1$Aghanim et al. (2020b); Escamilla et al. (2024a). [^5] In recent years, the possibility that the DE EoS can be phantom in nature has gained substantial interest, as in principle a shift of $\sim 20\%$ towards $w_{0}<-1$ could already be enough to address the well-known Hubble tension Verde et al. (2019); Di Valentino et al. (2021b); Perivolaropoulos and Skara (2022); Abdalla et al. (2022); Hu and Wang (2023) – see, e.g., Refs. Vagnozzi (2020); Giarè (2024) for an overview, as well as for caveats surrounding this possibility. However, as extensively documented in the literature, this preference is not confirmed by independent CMB experiments such as ACT and SPT Aiola et al. (2020); Dutcher et al. (2021); Balkenhol et al. (2023); Giarè (2023), and – most importantly – it lacks consistent support from observations of the local universe. When combining CMB, BAO and SN data altogether, no convincing deviation from $w_{0}=-1$ is seen, possibly lending weight to the cosmological constant interpretation Escamilla et al. (2024a).
@@
However, one may argue that simplicity may not always be a prerogative of nature. Pushing this approach forward, we can relax the assumption of a constant EoS and consider models where $w(a)$ varies with the universe’s expansion – here and after known as Dynamical Dark Energy (DDE). This possibility, along with the various proposed physical realizations, has also undergone extensive testing; see, e.g., Refs. Mainini et al. (2003); Alam et al. (2004); Sola and Stefancic (2006); Antoniadis et al. (2007); Yang et al. (2021b); Zhao et al. (2012); Solà Peracaula et al. (2018); Sola et al. (2017); Zhao et al. (2017); Yang et al. (2017, 2019a, 2019b); Sola Peracaula et al. (2019); Pan et al. (2019); Colgáin et al. (2021); Escamilla-Rivera and Nájera (2022); Zhao et al. (2020). From an observational standpoint, CMB data alone have limited capacity to constrain DDE models due to minimal effects left at the epoch of the last scattering surface and the increased number of cosmological parameters Di Valentino et al. (2017, 2020d, 2022). [^6] This difficulty is often referred to as geometrical degeneracy. At its core, the problem is that different combinations of late-time cosmic parameters can be adjusted in such a way that the acoustic angular scale $\theta_{s}$ – determined by the ratio of the comoving sound horizon at recombination to the comoving distance to last scattering – remains constant if both quantities change proportionally. Consequently, measurements based solely on this scale cannot provide strong constraints on (dynamical) DE parameters by themselves, unless perturbation-level effects and late-time data are also incorporated to break this degeneracy. Even in simple parametrizations aimed at minimizing the number of free degrees of freedom, CMB experiments produce constraints that are typically too broad to provide informative results. Therefore, local universe observations acquire primary importance.
A significant turning point in the study of DDE models is marked by the very recent BAO released by the Dark Energy Spectroscopic Instrument (DESI) Adame et al. (2024a); Lodha et al. (2024) and – albeit to a lesser extent – by SN distance moduli measurements resulting from the five-year observations of the Dark Energy Survey (DESY5) Abbott et al. (2024); Sánchez et al. (2024); Vincenzi et al. (2024). Before these updated data, no significant preference had ever emerged in favor of DDE models, certainly not to the extent required to challenge the baseline cosmological constant interpretation Gariazzo et al. (2024). Excitingly, when DESI BAO observations are combined with Planck CMB data and SN distance moduli measurements (either from the Pantheon-plus catalog Scolnic et al. (2022); Brout et al. (2022) or from DESY5 Abbott et al. (2024); Sánchez et al. (2024); Vincenzi et al. (2024)), they produce strong indications for DDE. Specifically, within the linear Chevallier-Polarski-Linder (CPL) parameterization of the DE EoS $w(a)=w_{0}+w_{a}(1-a)$ Chevallier and Polarski (2001); Linder (2003) – where $w_{0}$ represents the present-day DE EoS and $w_{a}$ quantifies the dynamical evolution – we observe a preference for $w_{0}>-1$ and $w_{a}<0$ at a statistical level ranging between 2.5 and 3.9 $\sigma$, depending on the specific combination of data involved Adame et al. (2024b).
@@
Unexpectedly, these results have heated up the recent debate, fueling a multitude of re-analyses and re-interpretations Cortês and Liddle (2024); Patel and Amendola (2024); Orchard and Cárdenas (2024); Liu et al. (2024); Chudaykin and Kunz (2024); Notari et al. (2024); Gialamas et al. (2024); Wang et al. (2024b); Lodha et al. (2024); Wang and Piao (2024); Giarè et al. (2024b); Carloni et al. (2024); Colgáin et al. (2024); Tada and Terada (2024); Yin (2024); Luongo and Muccino (2024); Park et al. (2024); Wang et al. (2024a); Shlivko and Steinhardt (2024); Ye et al. (2024); Li et al. (2024). As “extraordinary claims require extraordinary evidence”, we certainly advise and exercise caution. However, even taking in mind all the possible caveats and limitations surrounding the first DESI data release, it is undeniable that a high statistical preference for DDE holds intrinsic interest – if confirmed, this would represent the first concrete evidence of new physics beyond the standard model of cosmology.
@@
Given the potential implications of this result, it is natural to question its robustness. Barring any possible systematic issues in these datasets, we examine if and to what extent assuming a linear parameterization for the DE EoS might impact the current findings. Although the CPL parameterization has been demonstrated to match the background evolution of distances arising from exact DE equations of motion to an accuracy of approximately 0.1% for viable cosmologies over a wide range of physics, including scalar fields, modified gravity, and phase transitions, other parameterizations proposed over the years (which may deviate from CPL at both $z\ll 1$ and $z\gtrsim 1$) remain allowed by current observations. Testing these alternative models against new data can certainly represent a useful exercise to shed light on the role played by the parameterization itself. To be fair, the process has already begun with several independent groups actively engaged in this activity Hernández-Almada et al. (2024); Pourojaghi et al. (2024); Ramadan et al. (2024); Berghaus et al. (2024); Qu et al. (2024b); Notari et al. (2024); Adolf et al. (2024).
@@
Given the vast number of parameterizations proposed over the years and recently analyzed in relation to the DESI data, a few warnings are in order. First and foremost, alternative parameterizations often introduce extra parameters compared to CPL. This is both a blessing and a curse: on the one hand, accounting for more degrees of freedom allows more flexibility in $w(a)$. On the other hand, this typically implies relaxing the overall constraining power due to the combined effects of degeneracies and correlations among parameters. Secondly, the physical interpretations of the parameters involved may differ from the two employed in the CPL model. This further compounds the comparison of the results, making it difficult to derive general guidelines on the preference toward DDE.
@@
To overcome these difficulties, in this article, we test different DDE models while allowing for a fair comparison of the results. We restrict our attention to five well-known parameterizations that satisfy the following criteria: (i) they consist of two parameters to describe the evolution of $w(a)$, $w_{0}$ and $w_{a}$; (ii) these two parameters retain the same physical meaning as in the CPL parameterization; (iii) for the same combinations of pairs $w_{0}$ $w_{a}$, the resulting shape of $w(a)$ deviates from CPL either either near the present epoch or in the past, depending on the specific case.
@@
The paper is organized as follows. In Sec. II we sketch the theoretical set-up of the gravitational equations and propose the DE parametrizations we wish to study. In Sec. III we describe the observational data and the methodology to constrain the proposed DE parametrizations. In Sec. IV we present the constraints on the resulting DE scenarios. Finally, in Sec. V we draw our general conclusions.
II Dynamical Dark Energy Models

### v2

@@ Abstract
1 Introduction
One of the most undoubtedly fascinating and unforeseen discoveries of the past three decades is that the Universe is undergoing an accelerated phase of expansion. This was first argued in 1998 through observations of distant Type Ia Supernovae [1, 2], and has since been corroborated by a wide variety of other probes [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 45, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62]. [^2] For a few caveats, objections, and discussions surrounding this conclusion raised over the years, see Refs. [63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74].
@@ 1 Introduction
Since all known forces and components in nature would decelerate the expansion rate of the Universe, acceleration itself requires a physical mechanism beyond the Standard Model of fundamental interactions, able to counteract deceleration, inducing instead a Dark Energy (DE) phase where the dynamics is characterized by negative pressure with an effective Equation of State (EoS) $w<-1/3$, a condition that directly follows from the second Friedmann equation (also known as the acceleration equation) [75, 76, 77, 78, 79, 80, 81, 82].
@@
In the theoretical framework described by the standard $\Lambda$CDM model of cosmology, DE is parametrized by a positive cosmological constant term ($\Lambda$) in the Einstein equation with its energy density comprising the vast majority of the Universe’s energy budget. Despite its apparent simplicity, this interpretation is not free from conceptual problems and limitations [83, 84, 75, 85, 78, 86, 87, 88, 89, 90, 91, 92]. Foremost, plugging a positive cosmological constant component into the gravitational field equations described by General Relativity (GR) implies living in an asymptotically de Sitter universe, which seems to contrast with several theories/models of quantum gravity proposing instead an asymptotically anti-de Sitter universe [93, 94, 95, 96, 97]. Secondly, it seems quite natural to question what made us so lucky to live precisely in the cosmic epoch when such a constant component came to be not only relevant but even dominant compared to other contributions such as matter, altering the expansion history of the Universe so prominently as to allow us to become easily aware of its existence and implications [98, 99, 82, 100]. Finally, and most importantly, when it comes to the physical interpretations, anything that contributes to the energy density of the vacuum behaves akin to a cosmological constant, summing up within the energy-momentum tensor as $T_{\mu\nu}\propto g_{\mu\nu}\rho$ due to Lorentz invariance. Based on standard quantum field theory calculations, one would expect a zero-point energy density contribution of $\rho_{\rm vac}$, which, depending on the ultraviolet wavelength cutoff scale, is found to be somewhere between $10^{50}$ to $10^{120}$ orders of magnitude larger than what is inferred by cosmological data [83]. This leads to one of the biggest disagreements between theoretical predictions and observations, arguably requiring a level of fine-tuning that appears to be well beyond what any current theories can realistically explain [101, 83, 102, 78, 103]. As a result, from a theoretical perspective, the nature of DE remains one of the biggest puzzles in modern physics, sustaining significant research interest in the high-energy physics community. [^3] This diffuse interest is reflected in the wide range of models – both within and beyond the standard cosmological constant – that have been proposed over the years. These include, for example, new (ultra)light fields and modifications to gravity. With no claims to completeness see, e.g., Refs. [104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 92, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159, 160, 161, 162, 163, 164, 165, 166, 167].
@@
Despite these remarkable achievements, it is no exaggeration to say that at the time of writing, observations are inconclusive about the physical nature of DE. While too far-away deviations from the canonical $\Lambda$CDM model appear severely constrained [292], several alternative theoretical frameworks and phenomenological avenues featuring new physics in the dark sector of the model remain at the very least plausible.
@@
Trying to summarize an otherwise very articulated debate, we can adhere to Occam’s razor principle and start considering one of the simplest hypotheses beyond the cosmological constant. This involves assuming that DE can be modeled as a generic fluid with a constant EoS, $w_{0}$. By leaving $w_{0}$ as a free parameter in the theoretical model, cosmological observations can constrain deviations from the cosmological constant value ($w_{0}=-1$). [^6] See, e.g., Refs [293, 294, 295, 296, 297, 32, 298, 299, 300, 301, 302, 38, 303, 304, 305, 306, 307, 40, 308, 309, 45, 310, 311, 312, 313, 314, 54, 50, 315, 316, 317, 318, 319, 320, 321, 322, 323, 215, 324, 53, 325, 326, 327, 328, 329, 330, 331, 245, 332, 333, 334, 335, 336] for recent and not-so-recent discussions and constraints on the DE EoS from a variety of astrophysical and cosmological probes. In this regard, focusing exclusively on the Planck CMB data one might speculate about a preference for a phantom-like DE component ($w_{0}<-1$[40, 245]. [^7] In recent years, the possibility that the DE EoS can be phantom in nature has gained substantial interest, as in principle a shift of $\sim 20\%$ towards $w_{0}<-1$ could already be enough to address the well-known Hubble tension [337, 289, 290, 291, 338] – see, e.g., Refs. [197, 339] for an overview, as well as for caveats surrounding this possibility. However, as extensively documented in the literature, this preference is not confirmed by independent CMB experiments such as ACT and SPT [44, 51, 271, 340], and – most importantly – it lacks consistent support from observations of the local universe. When combining CMB, BAO and SN data altogether, no convincing deviation from $w_{0}=-1$ is seen, possibly lending weight to the cosmological constant interpretation [245].
@@
However, one may argue that simplicity may not always be a prerogative of nature. Pushing this approach forward, we can relax the assumption of a constant EoS and consider models where $w(a)$ varies with the universe’s expansion – here and after known as Dynamical Dark Energy (DDE). This possibility, along with the various proposed physical realizations, has also undergone extensive testing; see, e.g., Refs. [341, 342, 343, 344, 313, 345, 346, 347, 348, 349, 194, 195, 350, 351, 319, 352, 353]. From an observational standpoint, CMB data alone have limited capacity to constrain DDE models due to minimal effects left at the epoch of the last scattering surface and the increased number of cosmological parameters [354, 355, 356]. [^8] This difficulty is often referred to as geometrical degeneracy. At its core, the problem is that different combinations of late-time cosmic parameters can be adjusted in such a way that the acoustic angular scale $\theta_{s}$ – determined by the ratio of the comoving sound horizon at recombination to the comoving distance to last scattering – remains constant if both quantities change proportionally. Consequently, measurements based solely on this scale cannot provide strong constraints on (dynamical) DE parameters by themselves, unless perturbation-level effects and late-time data are also incorporated to break this degeneracy. Even in simple parametrizations aimed at minimizing the number of free degrees of freedom, CMB experiments produce constraints that are typically too broad to provide informative results. Therefore, local universe observations acquire primary importance.
[…]
@@
Unexpectedly, these results have heated up the recent debate, fueling a multitude of re-analyses and re-interpretations [359, 360, 361, 362, 363, 364, 365, 366, 59, 367, 253, 368, 369, 370, 371, 372, 373, 264, 374, 375, 266, 267]. As “extraordinary claims require extraordinary evidence”, we certainly advise and exercise caution. However, even taking in mind all the possible caveats and limitations surrounding the first DESI data release, it is undeniable that a high statistical preference for DDE holds intrinsic interest – if confirmed, this would represent the first concrete evidence of new physics beyond the standard model of cosmology.
@@
Given the potential implications of this result, it is natural to question its robustness. Barring any possible systematic issues in these datasets [^9] As argued in various recent works, the DESI BAO measurement at $z=0.71$ (which is in $\sim 3\sigma$ tension with Planck) can play a crucial role in deriving many of the DESI signals for new physics, partially including the preference for DDE [376, 369, 377]. , we examine if and to what extent assuming a linear parameterization for the DE EoS might impact the current findings. Although the CPL parameterization has been demonstrated to match the background evolution of distances arising from exact DE equations of motion to an accuracy of approximately 0.1% for viable cosmologies over a wide range of physics, including scalar fields, modified gravity, and phase transitions (see, e.g., Refs. [172, 378]), other parameterizations proposed over the years (which may deviate from CPL at both $z\ll 1$ and $z\gtrsim 1$) remain allowed by current observations. Testing these alternative models against new data can certainly represent a useful exercise to shed light on the role played by the parameterization itself. To be fair, the process has already begun with several independent groups actively engaged in this activity [379, 380, 381, 382, 383, 364, 384].
@@
Given the vast number of parameterizations proposed over the years and recently analyzed in relation to the DESI data, a few warnings are in order. First and foremost, alternative parameterizations often introduce extra parameters compared to CPL. This is both a blessing and a curse: on the one hand, accounting for more degrees of freedom allows more flexibility in $w(a)$. On the other hand, this typically implies relaxing the overall constraining power due to the combined effects of degeneracies and correlations among parameters. Secondly, the physical interpretations of the parameters involved may differ from the two employed in the CPL model. This further compounds the comparison of the results, making it difficult to derive general guidelines on the preference towards DDE.
@@
To overcome these difficulties, in this article, we test different DDE models while allowing for a fair comparison of the results. We restrict our attention to five well-known parameterizations that satisfy the following criteria: (i) they consist of two parameters to describe the evolution of $w(a)$, $w_{0}$ and $w_{a}$; (ii) these two parameters retain the same physical meaning as in the CPL parameterization; (iii) for the same combinations of pairs $w_{0}$ $w_{a}$, the resulting shape of $w(a)$ deviates from CPL either near the present epoch or in the past, depending on the specific case.
@@
The paper is organized as follows. In Sec. 2 we sketch the theoretical set-up of the gravitational equations and propose the DE parametrizations we wish to study. In Sec. 3 we describe the observational data and the methodology to constrain the proposed DE parametrizations. In Sec. 4 we present the constraints on the resulting DE scenarios. Finally, in Sec. 5 we draw our general conclusions.
2 Dynamical Dark Energy Models